OpenStudy (anonymous):

At Russell's school, one person will contact 4 people and each of those people will contact 4 other people, and so on. How many people will be contacted in the fifth round?

OpenStudy (anonymous):

First round = 4 people will be contacted. Second round = 4^2 (each person will contact four other people). Third round = 4^3 Forth Round = 4^4 Fifth Round = 4^5 = 1024.

OpenStudy (anonymous):

thx :)